OpenStudy (anonymous):

g(x) = 2(x − 4), find the value of x if g(x) = 20.

OpenStudy (anonymous):

20=2(x-4) divide both sides by 2: 10=x-4 add 4 to both sides: 14=x

OpenStudy (anonymous):

when i did it i got you add 8 to both side i know ur right but i wanna understand how to do these problems are you suppose to times the 2 by the -4 cuz thats what i did thats how i got -8

OpenStudy (anonymous):

well, i think what you did was this: